two firms in the same industry sell their product at 20$ per unit but one firm has tfc=200$ and avc=12$ while the other has tfc=600$ and avc's=6.66$
a. determine the breakeven output of each firm. why is the breakeven output of the second firm larger than that of the first firm ?
b. find the degree of the contribution margin for each firm at profit= 1200 for the first firm and at profit =1400 for the second firm
Answer #1

A.

Break even output for first firm = FC/(selling price - avc) = 200/(20-12)

Break even output for first firm = 25 units

Break even output for second firm = 600/(20-6.66)

Break even output for second firm = 44.98 or 45 units

Break even output for the second firm is larger than that of first firm, because second firm is having bigger fixed cost. Hence, second firm has to sell higher volume of products to recover the fixed cost.

---

B.

When profit for the first firm is $1200,

1200 = Contribution margin*25 + 200

Contribution margin = 1000/25

Contribution margin = $40

When profit for the second firm is $1400,

1400 = Contribution margin*45 + 600

Contribution margin = (1400-600)/45

Contribution margin = $17.78